OpenStudy (anonymous):

Need help with Linear Functions Unit Test 1.For the data in the table, does y vary directly with x? If it does write an equation for the direct varation. X Y 8 11 16 22 24 33 A.yes;y=2.75x B.yes;y=0.6875x C.yes;y=1.375x D.no; y does not vary directly with x 2.For the data in the table, does y vary directly with x? If it does, write an equation for the direct variation. X Y 40 32 28 16 16 12 A.yes;y=2x B.yes;y=0.5x C.yes;y=1.5x D.no; y does not vary directly with x

OpenStudy (bibby):

c d

OpenStudy (anonymous):

5.The table shows the number of miles driven over time. Time(hours) Distance(miles) 4 184 6 276 8 368 10 460 Express the relationship between distance and time in simplified form as a unit rate. Determine which statement correctly interprets this relationship. A.184; your car travels 184 miles. B.46/1; your care travels 16 miles in 1 hour. C.1/46; your car travels 46 miles in 1 hour D.4; your car travels 4 hours 7.What is the slope of the line that passes through the pair of points (2,5) and (8,3) A.1/3 B.-1/3 C.3 D.-3 8.What is the slope of the line that passes through the pair of points (3/2,-2) and (-3,7/3) A.-27/26 B.-26/27 C.26/27 D.27/26 10.Write an equation in point-slope form for the line through the given point with the given slope. (5,2); m=3 A.y+2=3(x-5) B.y+2=3x-5 C.y-2=3(x+5 D.y-2=3(x-5) 11.Write an equation in point-slope form for the line through the given point with the given slope. (-10,-6); m=-5/8 A.y-6=-5/8(x-10) B.y-6=-5/8(x+10) C.y+6=-5/8(x+10) D.y+6=-5/8(x+6) 12. The table shows the height of a plant as it grows. Which equation in point-slope form gives the plant's height at any time? Time(months) Plant height(cm) 2 18 4 36 6 54 8 72 A.y+18=9(x+2) B.y-18=9x+2 C.y-18=9(x-2) D.The relationship is nonlinear. 14.Write y=-2/3x+7 in standard form using integers. A.2x+3y=21 B.3x-2y=21 C.-2x-3y=21 D.-2x+3y=7 15.Write an equation for the line that is parallel to the given line and passes through the given point. y=3x+7; (2,10) A.y=3x+11 B.y=3x+4 C.-3x+4 D.y=(-1/3)x+4 16. Tell whether the lines for each pair of equations are parallel, perpendicular, or neither. y=-2x+4 -5x+10y=5 A.parallel B.perpendicular C.neither 17.Tell whether the lines for each pair of equations are parallel, perpendicular, or neither y=-1/4x+10 -2x+8y=6 A.parallel B.perpendicular C.neither 18.Write the equation of a line that is perpendicular to the given line and that passes through the given point. y-2=7/3(x+5); (-4,9) A.y-9=-3/7 (x-4) B.y-9=3/7 (x-4) C.y-9=-3/7 (x+4) D.y-9=3/7 (x+4) 20.Describe how the graphs y=|x| and y=|x+2| are related. A.The graphs have the same shape. The y-intercepts of y=|x| is 0, and the y-intercept of the second graph is -2 B.The graphs have the same shape.The y-intercepts of y=|x| is 0, and the x-intercept of the second graph is 2. C.The graphs have the same shape.The y-intercepts of y=|x| is 0, and the y-intercept of the second graph is 2. D.The two graphs are the same.

OpenStudy (iloveopenstudy2):

@ganeshie8 he's trying to get help on a test

OpenStudy (anonymous):

@bibby